# X-CrossNet: A complex spectral mapping approach to target speaker extraction with cross attention speaker embedding fusion

### v1

@@ Abstract
Target speaker extraction (TSE) is a technique for isolating a target speaker’s voice from mixed speech using auxiliary features associated with the target speaker. This approach addresses the cocktail party problem and is generally considered more promising for practical applications than conventional speech separation methods. Although academic research in this area has achieved high accuracy and evaluation scores on public datasets, most models exhibit significantly reduced performance in real-world noisy or reverberant conditions. To address this limitation, we propose a novel TSE model, X-CrossNet, which leverages CrossNet as its backbone. CrossNet is a speech separation network specifically optimized for challenging noisy and reverberant environments, achieving state-of-the-art performance in tasks such as speaker separation under these conditions. Additionally, to enhance the network’s ability to capture and utilize auxiliary features of the target speaker, we integrate a Cross-Attention mechanism into the global multi-head self-attention (GMHSA) module within each CrossNet block. This facilitates more effective integration of target speaker features with mixed speech features. Experimental results show that our method performs superior separation on the WSJ0-2mix and WHAMR! datasets, demonstrating strong robustness and stability.
@@ I Introduction
Blind source separation aims to separate all speaker channels from mixed speech signals. While BSS performs well on benchmark datasets like WSJ0-2mix [1], its application in real-world scenarios is challenging due to the unpredictable number of speakers. Target speaker extraction, which can be considered either an improvement on or a subfield of BSS, offers a more targeted approach. Unlike BSS, which relies solely on mixed speech as input, TSE also uses a reference speech segment from the target speaker. By leveraging the features of the reference speech, TSE identifies and extracts the target speaker while ignoring other speakers and background noise. From a technical perspective, TSE simplifies the complexity of BSS, and from an application standpoint, it is better suited to practical scenarios.
@@ III-C Speaker Extractor
The backbone network of the speaker extractor is derived from [15]. Similarly, we have introduced a random chunk positional encoding (RCPE) module after the speech encoder to generalize the model to longer sequences. As shown in Figure 2 (c), CrossNet mainly consists of $\mathrm{B}$-layer CrossNet Blocks, with a typical value of $\mathrm{B}$ is 12. Each block contains three main modules, global multi-head self-attention module (GMHSA), cross-band module, and narrow-band module. To fuse with speaker features ${\mathbf{E}}_{a}$, we have improved the GMHSA of each block layer by adding a cross-attention-based fusion structure as shown in Figure 2 (d). The cross-band module consists of two frequency convolutional modules and a full-band linear module, where the frequency convolutional module contains a layer normalization (LN) layer, A grouped convolution layer along the frequency axis, and a parametric rectified linear unit (PReLU) activation function. The full-band linear module includes a linear layer, a sigmoid-weighted linear unit (SiLU), and a set of linear layers along the frequency axis The narrow-band module is a modified version of the Conformer [16] convolutional block, which includes an LN layer, a linear layer, a SiLU activation, a time-convolutional (T-Conv) layer and a final linear layer.
@@ III-E Loss functions
${\mathcal{L}}_{\mathrm{SI-SDR}}=-10\log_{10}{\frac{{\left\|{\alpha}{s}_{1}\right\|}_{2}^{2}}{{\left\|\hat{{s}}-{\alpha}{s}_{1}\right\|}_{2}^{2}}}$ (7)
${\alpha}=\frac{\left\langle\hat{s},{s}_{1}\right\rangle}{\left\langle{s}_{1},{s}_{1}\right\rangle}$ (8)
@@ IV-C Evaluation results
We use signal-to-distortion ratio improvement (SDRi) [26] and SI-SDR improvement (SI-SDRi) [17] as evaluation metrics, where higher values indicate better model performance. First, we compared the performance of X-CrossNet with state-of-the-art (SOTA) models on the WSJ0-2mix dataset. The results are presented in Table I. The second column of the table specifies the input type for each method: X-TF GridNet and X-CrossNet operate in the time-frequency domain, while the other methods function in the time domain. As shown in the table, X-CrossNet outperforms other methods on both SDRi and SI-SDRi metrics and achieves this with the lowest number of parameters. For experiments on the WHAMR! dataset, which includes noise and reverberation conditions, results are presented in Table II. The third column of the table shows the length of the reference speech used, with 7.3 seconds selected for comparison. The results demonstrate that the proposed method achieves superior performance compared to existing approaches, confirming that X-CrossNet exhibits strong robustness and stability in noisy and reverberant scenarios.
@@ V Conclusion
In this paper, we propose a time-frequency domain target speaker extraction method named X-CrossNet, which uses CrossNet as the backbone network and is designed to perform effectively in noisy and reverberant environments. We introduce a REL block based speaker encoder to extract accurate speaker representations and enhance the GMHSA module in CrossNet by incorporating a fusion structure for speaker embeddings. Experimental results demonstrate that X-CrossNet achieves state-of-the-art performance on conventional test sets while maintaining significant robustness and stability on noisy and reverberant datasets.

### v2

@@ Abstract
Target speaker extraction (TSE) is a technique for isolating a target speaker’s voice from mixed speech using auxiliary features associated with the target speaker. It is another attempt at addressing the cocktail party problem and is generally considered to have more practical application prospects than traditional speech separation methods. Although academic research in this area has achieved high performance and evaluation scores on public datasets, most models exhibit significantly reduced performance in real-world noisy or reverberant conditions. To address this limitation, we propose a novel TSE model, X-CrossNet, which leverages CrossNet as its backbone. CrossNet is a speech separation network specifically optimized for challenging noisy and reverberant environments, achieving state-of-the-art performance in tasks such as speaker separation under these conditions. Additionally, to enhance the network’s ability to capture and utilize auxiliary features of the target speaker, we integrate a Cross-Attention mechanism into the global multi-head self-attention (GMHSA) module within each CrossNet block. This facilitates more effective integration of target speaker features with mixed speech features. Experimental results show that our method performs superior separation on the WSJ0-2mix and WHAMR! datasets, demonstrating strong robustness and stability.
@@ I Introduction
Blind source separation aims to separate all speaker channels from mixed speech signals. While BSS performs well on benchmark datasets like WSJ0-2mix [1], its application in real-world scenarios is challenging due to the unpredictable number of speakers. Target speaker extraction, which can be considered either an improvement on or a subfield of BSS, offers a more feasible approach. Unlike BSS, which relies solely on mixed speech as input, TSE also uses a reference speech segment from the target speaker. By leveraging the features of the reference speech, TSE identifies and extracts the target speaker while ignoring other speakers and background noise. From a technical perspective, TSE simplifies the complexity of BSS, and from an application standpoint, it is better suited to practical scenarios.
@@ III-C Speaker Extractor
The backbone network of the speaker extractor is derived from [15]. Similarly, we have introduced a random chunk positional encoding (RCPE) module after the speech encoder to generalize the model to longer sequences. As shown in Figure 2 (c), CrossNet mainly consists of $\mathrm{B}$-layer CrossNet Blocks. Each block contains three main modules, global multi-head self-attention module (GMHSA), cross-band module, and narrow-band module. To fuse with speaker features ${\mathbf{E}}_{a}$, we have improved the GMHSA of each block layer by adding a cross-attention-based fusion structure as shown in Figure 2 (d). The cross-band module consists of two frequency convolutional modules and a full-band linear module, where the frequency convolutional module contains a layer normalization (LN) layer, A grouped convolution layer along the frequency axis, and a parametric rectified linear unit (PReLU) activation function. The full-band linear module includes a linear layer, a sigmoid-weighted linear unit (SiLU), and a set of linear layers along the frequency axis The narrow-band module is a modified version of the Conformer [16] convolutional block, which includes an LN layer, a linear layer, a SiLU activation, a time-convolutional (T-Conv) layer and a final linear layer.
@@ III-E Loss functions
${\mathcal{L}}_{\mathrm{SI-SDR}}=-10\log_{10}{\frac{{\left\|{\alpha}{s}_{1}\right\|}_{2}^{2}}{{\left\|\hat{{s}}-{\alpha}{s}_{1}\right\|}_{2}^{2}}},{\alpha}=\frac{\left\langle\hat{s},{s}_{1}\right\rangle}{\left\langle{s}_{1},{s}_{1}\right\rangle}$ (7)
${\mathcal{L}}_{\mathrm{CE}}=-\displaystyle\sum_{n=1}^{{N}_{s}}{p}_{n}\log\left({{\hat{p}}_{n}}\right)$ (8)
@@ IV-C Evaluation results
We use signal-to-distortion ratio improvement (SDRi) [26] and SI-SDR improvement (SI-SDRi) [17] as evaluation metrics, where higher values indicate better model performance. First, we compared the performance of X-CrossNet with state-of-the-art (SOTA) models on the WSJ0-2mix dataset. The results are presented in Table I. The second column of the table specifies the input type for each method: X-TF-GridNet and X-CrossNet operate in the time-frequency domain, while the other methods function in the time domain. As shown in the table, X-CrossNet outperforms other methods on both SDRi and SI-SDRi metrics and achieves this with the lowest number of parameters. For experiments on the WHAMR! dataset, which includes noise and reverberation conditions, results are presented in Table II. The third column of the table shows the length of the reference speech used, with 7.3 seconds selected for comparison. The results demonstrate that the proposed method achieves superior performance compared to existing approaches, confirming that X-CrossNet exhibits strong robustness and stability in noisy and reverberant scenarios.
@@ V Conclusion
In this paper, we propose a time-frequency domain target speaker extraction method named X-CrossNet, which uses CrossNet as the backbone network and is designed to perform effectively in noisy and reverberant environments. We introduce a REL block based speaker encoder to extract accurate speaker representations and enhance the GMHSA module in CrossNet by incorporating a fusion structure for speaker embeddings. Experimental results demonstrate that X-CrossNet achieves state-of-the-art performance on conventional test sets while maintaining significant robustness and stability in noisy and reverberant scenarios.